Gucchi:

Simplify the following expression: 4n − m + 3p + 2m − n − p 3n + m + 2p 3n + 5m − 2p 5n + 3m + 4p 5n + m − p

Gucchi:

Am I right with A?

Gucchi:

this is easy

stopscammingpeople:

i wana help but i feel like ima get it wrong lolll .

stopscammingpeople:

but lets break it down

stopscammingpeople:

you're simplying the expression.. right?

Gucchi:

i mean it says that in the problem

stopscammingpeople:

it was yes or no lol ok anyway.

stopscammingpeople:

did you work out the problem yet?? or do i have to help you w that too?

Gucchi:

i did

stopscammingpeople:

ok so , since you've already did it .. we can cross out b . we know b is wrong for sure .

carmelle:

@stopscammingpeople you want me to help or u got this?

Gucchi:

B and c are crossed out

carmelle:

@stopscammingpeople wrote:
ok so , since you've already did it .. we can cross out b . we know b is wrong for sure .
you cant jus say that tho- u gotta tell him why B is wrong

stopscammingpeople:

i know the answer its just a lil hard to explain .

Gucchi:

you wouldnt get 5n cause subtraction

stopscammingpeople:

yea you're right .

@carmelle wrote:
@stopscammingpeople wrote:
ok so , since you've already did it .. we can cross out b . we know b is wrong for sure .
you cant jus say that tho- u gotta tell him why B is wrong

Gucchi:

a is right so yh

stopscammingpeople:

you just answered your own question ..

Gucchi:

m is by itself no nummber

stopscammingpeople:

bro are you sure you need help w these or are you just using random people to ensure that youre right?

mercades74605:

fr

snowflake0531:

@gucchi wrote:
Am I right with A?
Yes, you are correct

stopscammingpeople:

you cant tell him that

carmelle:

lemme help, we are basically combining like terms, lets start with `n` `4n` − m + 3p + 2m `− n` − p we know that when the is no number in front of a variable, it is equal to 1 so, 4n - 1n = ?

stopscammingpeople:

carm this guy is making my head hurt lol

Gucchi:

@carmelle wrote:
lemme help, we are basically combining like terms, lets start with `n` `4n` − m + 3p + 2m `− n` − p we know that when the is no number in front of a variable, it is equal to 1 so, 4n - 1n = ?
3n

stopscammingpeople:

bro aint worried about us anymore he just wanna make sure he get that A lol

Gucchi:

@stopscammingpeople wrote:
bro aint worried about us anymore he just wanna make sure he get that A lol
wouldnt anyone focus on getting an A?

carmelle:

right, so that means we can eliminate C and D.

Gucchi:

bro if you getting a headache dont click on my questions

stopscammingpeople:

bro .. its not the question

stopscammingpeople:

its you .

stopscammingpeople:

you had the answer before you came to qc

XxXNessalulbaddieXxX:

theres nothing aganist putting questions up to make sure ur right

stopscammingpeople:

@xxxnessalulbaddiexxx wrote:
theres nothing aganist putting questions up to make sure ur right
facts indeed.

stopscammingpeople:

but still

stopscammingpeople:

he said "ITs A sO"

stopscammingpeople:

As if i didnt just help him lol

snowflake0531:

you didn't

carmelle:

y'all needa chill, anyways lets continue here lets go to `m` now 4n `− m` + 3p `+ 2m` − n − p -m + 2m = ?

stopscammingpeople:

you're right lowkey

stopscammingpeople:

i was tryna help him get the answer

stopscammingpeople:

but you just told him anyway

XxXNessalulbaddieXxX:

@stopscammingpeople wrote:
you're right lowkey
just stop spamming the reply section focus on the question that was asked

stopscammingpeople:

sorry ok , lets get back on track

stopscammingpeople:

@carmelle wrote:
y'all needa chill, anyways lets continue here lets go to `m` now 4n `− m` + 3p `+ 2m` − n − p -m + 2m = ?
how do you put the math problem in a equation like that??

Gucchi:

@carmelle wrote:
y'all needa chill, anyways lets continue here lets go to `m` now 4n `− m` + 3p `+ 2m` − n − p -m + 2m = ?
3m

Gucchi:

wait

snowflake0531:

\(-m+2m=2m-m=2m-1m\) try again

carmelle:

no.. pay attention to the signs

Gucchi:

m?

carmelle:

yes thats correct, since we know that, which choice can we eliminate now?

Gucchi:

b & c

carmelle:

yup, and that leaves us with..?

mercades74605:

yall teachin me how to do this 2 cuz i suck at algebra thank you

Gucchi:

A

Gucchi:

and d

carmelle:

YESSSS

Gucchi:

but a is right

carmelle:

we already eliminated D 💀

Gucchi:

oh wait wat

Gucchi:

i didnt recognize lol

carmelle:

@carmelle wrote:
right, so that means we can eliminate C and D.
remember...

mercades74605:

in conclusion yes you were right with A

Gucchi:

but yh 3p - p is 2p

carmelle:

@gucchi wrote:
but yh 3p - p is 2p
yes

Gucchi:

@carmelle wrote:
@carmelle wrote:
right, so that means we can eliminate C and D.
remember...
ohhh my bad didnt pay attention

carmelle:

its alright :)